Iggy Company is considering three capital expenditure projects. Relevant data for the projects are as follows.

Project Investment Annual
Income
Life of
Project
22A $243,500 $17,150 6 years
23A 272,500 20,770 9 years
24A 284,200 15,700 7 years


Annual income is constant over the life of the project. Each project is expected to have zero salvage value at the end of the project. Iggy Company uses the straight-line method of depreciation.

Determine the internal rate of return for each project. (Round answers 0 decimal places, e.g. 10. For calculation purposes, use 5 decimal places as displayed in the factor table provided.)
Answer #1

Depreciation = (cost - salvage value) / life

22A = 243500/6 = 40583

23A = 272500/9 = 30278

24A = 284200/7 = 40600

Annual cash inflows = net income + depreciation

22A = 17150+40583 = 57733

23A = 20770+30278 = 51048

24A = 15700+40600 = 56300

Internal rate of return = initial investment / annual cash inflows = found the value in present value annuity $1 table

22A = 243500 /57733 = 11%

23A = 272500/51048 = 12%

24A = 284200 / 56300 = 9%